OpenStudy (anonymous):

Graph each function. Identify the domain and range. -3x if x <_ -4 f(x)= x if 0 < x <_ 3 8 if x > 3 Help!!!

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

How to answer these?

OpenStudy (anonymous):

I really am desperate to learn this coz i really dnt know anything about graphing

jimthompson5910 (jim_thompson5910):

this is a piecewise function, so you graph each piece to graph the overall piecewise function

OpenStudy (anonymous):

Yeah i dont know. :(

jimthompson5910 (jim_thompson5910):

how would you graph the first piece how would you graph y = -3x

OpenStudy (anonymous):

Thats what i dont know.

jimthompson5910 (jim_thompson5910):

when x = 0, what is y?

OpenStudy (anonymous):

Imreally need help on this one

jimthompson5910 (jim_thompson5910):

I am helping, y = -3x y = -3*0 ... plug in x = 0 y = ???

OpenStudy (anonymous):

How couild i plug to 0 oh i get it on the left side? Or down?

jimthompson5910 (jim_thompson5910):

look it how I replaced x with 0

jimthompson5910 (jim_thompson5910):

you then evaluate and simplify

OpenStudy (anonymous):

Ill identifynthe domain and range first then ill do graphing but first i really dont understand what are u talking about

jimthompson5910 (jim_thompson5910):

y = -3x y = -3*0 ... plug in x = 0 (ie replace x with 0) y = ???

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

-3 times 0 is equal to -3?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

yes, so when x = 0, y = 0

jimthompson5910 (jim_thompson5910):

when x = 1, what is y?

OpenStudy (anonymous):

1?

jimthompson5910 (jim_thompson5910):

y = -3x y = -3*1 ... plug in x = 1 (ie replace x with 1) y = ???

OpenStudy (anonymous):

-3 plug graph?

jimthompson5910 (jim_thompson5910):

when x = 2, y = ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

use the equation we've been using

OpenStudy (anonymous):

1?

jimthompson5910 (jim_thompson5910):

no, the equation is y = -3x and you replace or plug in x = 2

OpenStudy (anonymous):

Then

jimthompson5910 (jim_thompson5910):

plug it in and tell me what you get

OpenStudy (anonymous):

What do you mean what i get on the graph?

OpenStudy (anonymous):

Cnan u draw i really cant i need to see an explanation

jimthompson5910 (jim_thompson5910):

to graph y = -3x, plot the two points (0,0) and (1,-3) then draw a line through them

jimthompson5910 (jim_thompson5910):

each point is found by plugging in a number for x to get a corresponding number for y

OpenStudy (anonymous):

I dont have an idea to graph

OpenStudy (anonymous):

But i think i got it just help me with the domain and range

jimthompson5910 (jim_thompson5910):

you're just plotting those points, then drawing a line through them

OpenStudy (anonymous):

How about domain and range

jimthompson5910 (jim_thompson5910):

what does your graph look like

OpenStudy (anonymous):

Just connecting the points and 0 to 0 just infnity

jimthompson5910 (jim_thompson5910):

can you draw it out to show me what you got?

OpenStudy (anonymous):

How will i draw a graph here?

OpenStudy (anonymous):

I still have like 14 questions to answer. :(

jimthompson5910 (jim_thompson5910):

use the draw button under the message box

OpenStudy (anonymous):

I'm using ipad i think there is no draw button.

jimthompson5910 (jim_thompson5910):

really? hmm how odd

jimthompson5910 (jim_thompson5910):

try scrolling down maybe?

OpenStudy (anonymous):

There is no draw button

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

well the domain and range is found by looking at the graph

jimthompson5910 (jim_thompson5910):

and determining which x values are allowed and which y values are possible

OpenStudy (anonymous):

0,0 -2,2?

OpenStudy (anonymous):

-3 i mean

jimthompson5910 (jim_thompson5910):

look at your graph, which x values are not allowed

OpenStudy (anonymous):

What do you mean are not allowed i dontnreally get it 😂

jimthompson5910 (jim_thompson5910):

which values can you not plug in the function

OpenStudy (anonymous):

What?

jimthompson5910 (jim_thompson5910):

what is the definition of domain

OpenStudy (anonymous):

All x values?

OpenStudy (anonymous):

Get it!

jimthompson5910 (jim_thompson5910):

what about the x values

OpenStudy (anonymous):

How can i write the piecewise function can undraw for me an example with an answer

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.mathsisfun.com/sets/domain-range-codomain.html and it explains what domain and range is

OpenStudy (anonymous):

Inreally dont have future woth his one ill let this pass for now

jimthompson5910 (jim_thompson5910):

have a look at that page a few times

OpenStudy (anonymous):

How about f(x)=[x]-6

jimthompson5910 (jim_thompson5910):

and anything you need graphed, you can use geogebra to help you do it

OpenStudy (anonymous):

I dnt have an idea how to answer that

jimthompson5910 (jim_thompson5910):

well you just need a bunch of points, plug in values of x to get output y values

jimthompson5910 (jim_thompson5910):

and is that supposed to be f(x) = |x|-6

OpenStudy (anonymous):

Okay what points should i plug?

jimthompson5910 (jim_thompson5910):

any x values you want (keep them small preferably)

OpenStudy (anonymous):

Whats my 6b

OpenStudy (anonymous):

Y i mean

jimthompson5910 (jim_thompson5910):

6b?

jimthompson5910 (jim_thompson5910):

not sure what you're asking

OpenStudy (anonymous):

Inhave x value whats my y value

jimthompson5910 (jim_thompson5910):

what's your x value

OpenStudy (anonymous):

0.5

jimthompson5910 (jim_thompson5910):

ok so replace x with that value then evaluate

OpenStudy (anonymous):

Ok nevermindni really cant do this. Nex time